OpenStudy (dehelloo):

find the value of x if determinant [3][2] = 9 [6][x]

OpenStudy (johnweldon1993):

Hint, determinant of a 2x2 matrix is \[Det\left[\begin{matrix}a & b \\ c & d\end{matrix}\right] = ad - bc\]

OpenStudy (dehelloo):

3 x X - 12 = 9

OpenStudy (dehelloo):

how do we determine X

OpenStudy (johnweldon1993):

Well if \(\large 3x - 12= 9\) How can we use algebra to isolate and solve for 'x'?

OpenStudy (dehelloo):

Do we divide 3 first to get it to the other side or subtact 12 from 9 first?

OpenStudy (johnweldon1993):

Here, you could do either So lets divide everything by 3 then \[\large \frac{3x - 12 = 9}{3} \rightarrow x - 4 = 3\] What now?

OpenStudy (dehelloo):

so then x just equals 7

OpenStudy (johnweldon1993):

Correct

OpenStudy (dehelloo):

So do we always begin with multiplying/dividing first?